# UniDream: Unifying Diffusion Priors for Relightable Text-to-3D Generation

### v1

@@ Abstract
Recent advancements in text-to-3D generation technology have significantly advanced the conversion of textual descriptions into imaginative well-geometrical and finely textured 3D objects. Despite these developments, a prevalent limitation arises from the use of RGB data in diffusion or reconstruction models, which often results in models with inherent lighting and shadows effects that detract from their realism, thereby limiting their usability in applications that demand accurate relighting capabilities. To bridge this gap, we present UniDream, a text-to-3D generation framework by incorporating unified diffusion priors. Our approach consists of three main components: (1) a dual-phase training process to get albedo-normal aligned multi-view diffusion and reconstruction models, (2) a progressive generation procedure for geometry and albedo-textures based on Score Distillation Sample (SDS) using the trained reconstruction and diffusion models, and (3) an innovative application of SDS for finalizing PBR generation while keeping a fixed albedo based on Stable Diffusion model. Extensive evaluations demonstrate that UniDream surpasses existing methods in generating 3D objects with clearer albedo textures, smoother surfaces, enhanced realism, and superior relighting capabilities.
{strip}
@@
* Equal contributions
@@ 1 Introduction
Recent methods [31, 16, 21, 4, 39] have been exploring the generation of 3D assets from textual descriptions under the supervision of 2D diffusion model [34, 33, 35]. For example, DreamFusion [31] defines a learnable Neural Radiance Fields (NeRF) [22] and optimizes it based on the Score Distillation Sampling (SDS). To enhance generation quality, subsequent studies have diversified the pipeline, focusing on aspects like 3D representations [16], loss functions [39], 3D prior [44], and 2D diffusion models [37, 47]. Although these methods achieve impressive results, they cannot generate relightable objects, as they typically represent the underlying illumination and texture of an object as a holistic appearance, as shown in the first four columns in Fig.1, which results in inherent lighting and shadows baked into the texture of the generated object. When relighting, the inherent highlights and shadows on these textured surfaces can affect the realism of the object.
In this paper, we present UniDream, a novel framework that allows generating relightable objects from textual descriptions. Fig.2 shows the fundamental difference between our method and other existing methods.
Our key idea is training a diffusion model that can provide both Physically-Based Rendering (PBR) material prior and multi-view geometry prior. Specifically, we first develop an albedo-normal aligned multi-view diffusion model (AN-MVM) for consistent multi-view image generation, which is trained on a set of paired albedo and normal data rendered from 3D object datasets. Then, following the simplified Diesney BRDF model [2], we define a 3D representation that includes albedo, normal, roughness, and metallic properties, which are optimized based on the trained diffusion model and Stable Diffusion [34] model. Compared with previous text-to-3D methods [31, 16, 39, 37], our approach is able to disentangle the illumination and PBR material, achieving high-quality relightable objects under different ambient lighting conditions, as shown in last three columns of Fig. 1.
@@ 2.1 Text-to-3D Generation
In recent years, the field of text-to-3D content generation has seen significant advancements, largely inspired by the advancements in text-to-image generation. These advances have particularly been driven by methods employing CLIP-based guidance or score distillation. Methods like  [12, 42, 23] utilize the text-image relationship inherent in CLIP [32] to facilitate general text-to-3D content creation. Another innovative approach, pioneered by DreamFusion [31], employs score distillation to enhance the robustness of text-to-3D generation, which also has been further developed and expanded in subsequent works [16, 39, 45]. The central to both of these paradigms is the use of pre-trained text-to-image diffusion models as a foundation, enabling the creation of diverse and imaginative 3D content. Furthermore, recent advancements have been developed by works [47, 37, 18], which innovatively employ models derived from the Objaverse [6] 3D dataset. These models are used to render 2D RGB images, aiming to train text-to-multi-view image diffusion models. This approach represents a significant stride in optimization-based 3D model generation, as it enables the simultaneous generation of multiple coherent images. Such a technique effectively addresses the challenge of inconsistent directions in geometric optimization, thereby enhancing the consistency and quality of 3D model generation.
In a different paradigm, some methods have shifted towards training diffusion or reconstruction models directly on paired text-3D data. This strategy enables the creation of 3D models that inherently possess text-like semantics. A variety of techniques [5, 48, 11, 27, 13, 8] fall under this category of 3D diffusion generation. In these models, textual information serves as a conditional input, guiding the generation process. This approach emphasizes the manipulation of the underlying 3D data representation, ensuring that the resulting models are both semantically rich and accurate representations of the text descriptions. Moreover, innovative strides have been taken in the realm of 3D reconstruction methods, particularly those grounded in transformer models, exemplified by LRM [10] and Instant3D [15]. These methods introduce a novel perspective, enabling the generation of high-quality 3D models from text or images within seconds, courtesy of their efficient reconstruction networks. The adoption of 3D diffusion and reconstruction methodologies has gained prominence due to their impressive speed in generating 3D objects.
The 2D multi-view diffusion approach and the 3D reconstruction technique utilizing 3D data have provided substantial inspiration. UniDream is constructed with integrating the strengths of both these methods.
@@ 2.2 Materials Generation
Drawing inspiration from recent material estimation techniques, newer research has focused on generating surface materials for 3D objects. For instance, Fantasia3D [3] combines a physical differential rendering pipeline with SDS to produce detailed 3D objects with realistic surface materials. However, this approach sometimes mixes albedo with reflected light, resulting in blurred material property distinctions. To address this, MATLABER [43] employs a latent BRDF auto-encoder, trained on an extensive dataset, to more effectively differentiate these properties. Building upon this, our method initiates with fixed albedo and normal generation and progressively incorporates other BRDF parameters, achieving a more natural and effective decomposition of surface materials in 3D objects.
@@ 3.1 Albedo-Normal Aligned Multi-view Consistent Diffusion Model
Multi-Domain Self-Attention. Based on multi-view consistency and sharing a similar perspective to recent work Wonder3D [19], we further introduce multi-domain consistency. Specifically, we introduce a distinct class label $L$ for the normal domain, and use a two-layer multi-layer perceptron (MLP) to encode this class label $L$ to obtain feature $F_{l}$ with the same dimension with the time-step features $F_{t}$, and $F_{l}$ is added to $F_{t}$ to control the generation process within the normal domain. Subsequently, we apply the self-attention mechanism to the corresponding views between the albedo and normal domains to ensure domain consistency.
@@
Text and Image Semantic Alignment. In order to solve the problem of potential semantic generalization loss caused by only using less 3D synthetic data during AN-MVM training, we use joint training to combine the 2D LAION-Aesthetics data with the 3D data we collected. Specifically, in line with the approach used by MVDream [37] during our training process, each batch is randomly composed by 3D data or 2D data based on a probability distribution: there’s a 70% chance of using 3D data and a 30% chance of incorporating 2D LAION-Aesthetics data. In this setting, while the normal domain is distinguished by class label, the albedo and 2D LAION-Aesthetics data are not. Moreover, to further differentiate between 3D and 2D data, we add the phrase ”, 3D asset” to the captions of 3D data. This strategic inclusion of a significant proportion of 2D data plays a crucial role in enhancing the semantic alignment between the generated image content and the corresponding input text.
@@ 3.2 Transformer-Based Reconstruction Model
Inspired by LRM [10] and Instant3D [15], we have integrated reconstruction models into our text-to-3D generation pipeline, aiming to provide an initial 3D prior for enhancing text-to-3D generation performance. As illustrated in the TRM module of Fig.3, for each object in the AN-MVM training dataset, we randomly select four views $I_{i}$, identical elevation but orthogonal views. Along with these views, the corresponding camera parameters, $C_{i}$, are used as inputs to the model. In this framework, the pre-trained ViT-based DINO-v2 [29] model, denoted as $F$ is employed to extract image features, $F_{i}$, from the albedo images of the four selected views. Concurrently, a learnable camera modulation module processes the camera parameters for each view using $\text{MLP}^{\text{mod}}(C_{i})$ and seamlessly integrates these encoded parameters into the image features. Then we employ learnable tokens, denoted as $T_{h}$, as input to the transformer model. These tokens are designed to undergo cross-attention with the image features, allowing for an effective fusion of the input image information. Subsequently, this is followed by the integration of cross-attention and self-attention modules and a multi-layer perceptron (MLP) to form a transformer block. Multiple such transformer blocks work in sequence to decode the input tokens $T_{h}$ into refined triplane representations $T_{h}^{\prime}$. The representations correspond to the semantic information of the input image. Finally, we decode these triplane representations using a MLP decoder of NeRF to reconstruct the final 3D model, denoted as $\text{Gen}_{\text{3D}}$. This entire process is detailed in Eq.1. $M$ is the number of transformer layers.
@@
Our approach differs from the configurations used in Instant3D in several key aspects to better adapt our model’s requirements. Firstly, we remove the intrinsic camera parameters of the input multi-view images and only normalize and encode the extrinsic parameters to adapt the multi-view images output by AN-MVM. Furthermore, we use albedo instead of RGB for training to prevent the impact of lighting and shadows in RGB images on the triplane-NeRF reconstruction results. Additionally, in order to increase the resolution of reconstructed model and save training cost, we resize the reference views between $128\times 128$ and $256\times 256$ resolution, and randomly crop $128\times 128$ images from the resized views to supervise the aligned region of rendered images. Finally, in the supervision, we not only render albedo for supervision but also incorporate normal supervision to significantly accelerate the model’s convergence speed and promote the geometry details. These strategic enhancements enable our TRM to deliver superior reconstruction results while reducing the training cost.
@@ 3.4 Physically-Based Rendering (PBR) Material Generation
Based on the geometry and albedo from the DMTet refinement, we employ the Stable Diffusion [34] model to generate the PBR material. We adopt the PBR material and illumination representations in Nvdiffrec [24], which is recognized for its speed and efficiency.
@@ 4.1 Implementation Details
Score Distillation Sample(SDS) Refining Details. We implement the refinement stages of NeRF and DMTet based on the Threestudio [^2] https://github.com/threestudio-project/threestudio. Specifically, we render four-view albedo and normal maps at the same time for SDS training. In the NeRF and DMTet refinement stages, we train 5,000 and 2,000 iterations respectively. During training, the loss weights of the albedo and normal domains are 0.8 and 0.2, and we use an ‘unsharp’ operation [40] is used in the last 500 iterations of each stage.
4.2 Qualitative Comparisons
Figure 5: Comparison of multi-view results generated by MVDream and UniDream.
Figure 6: 3D reconstruction performance of our UniDream.
4.3 Quantitative Evaluations
@@ 5 Ablation and Analysis
Multi-view Reconstruction Visualization. By employing multiple views (as depicted in Fig.5) as input, TRM reconstructs the corresponding 3D model, subsequently rendering albedo and normal maps from various views, as shown in Fig.6. The results demonstrate that our TRM successfully reconstructs 3D models closely resembling the input images in terms of geometry and texture semantics. This provides a robust prior for SDS refinement, thereby enhancing the stability and convergence speed of SDS-based training.
@@
Environment Maps. Fig.8 shows the different environment maps used by UniDream. Fig.8(a) is the environment map used to Fig.7 and the visualize of overall capabilities of UniDream. Fig.8(b) and Fig.8(c) are the environment maps used by ’Relighting-I’ and ’Relighting-II’ in the teaser of UniDream respectively.

### v2

@@ Abstract
Recent advancements in text-to-3D generation technology have significantly advanced the conversion of textual descriptions into imaginative well-geometrical and finely textured 3D objects. Despite these developments, a prevalent limitation arises from the use of RGB data in diffusion or reconstruction models, which often results in models with inherent lighting and shadows effects that detract from their realism, thereby limiting their usability in applications that demand accurate relighting capabilities. To bridge this gap, we present UniDream, a text-to-3D generation framework by incorporating unified diffusion priors. Our approach consists of three main components: (1) a dual-phase training process to get albedo-normal aligned multi-view diffusion and reconstruction models, (2) a progressive generation procedure for geometry and albedo-textures based on Score Distillation Sample (SDS) using the trained reconstruction and diffusion models, and (3) an innovative application of SDS for finalizing PBR generation while keeping a fixed albedo based on Stable Diffusion model. Extensive evaluations demonstrate that UniDream surpasses existing methods in generating 3D objects with clearer albedo textures, smoother surfaces, enhanced realism, and superior relighting capabilities. The project homepage is at: https://UniDream.github.io.
Keywords:
Text-to-3D Generation Multi-view Diffusion PBR Material
@@
* Equal $\dagger$ Corresponding author
@@ 1 Introduction
Recent methods [34, 18, 24, 4, 45, 55] have been exploring the generation of 3D assets from textual descriptions under the supervision of 2D diffusion models [37, 38, 60]. For example, DreamFusion [34] defines a learnable Neural Radiance Fields (NeRF) [25] and optimizes it based on the Score Distillation Sampling (SDS). To enhance generation quality, subsequent studies have diversified the pipeline, focusing on aspects like 3D representations [18], loss functions [45], 3D prior [54], and 2D diffusion models [40, 58]. Although these methods achieve impressive results, they cannot generate relightable objects, as they typically represent the underlying illumination and texture of an object as a holistic appearance, as shown in the first four columns in Fig.1, which results in inherent lighting and shadows baked into the texture of the generated object. When relighting, the inherent highlights and shadows on these textured surfaces can affect the realism of the object.
In this paper, we present UniDream, a novel framework that allows generating relightable objects from textual descriptions. Fig.2 shows the fundamental difference between our method and other existing methods. Our key idea is training a diffusion model that can provide both Physically-Based Rendering (PBR) material prior and multi-view geometry prior. Specifically, we first develop an albedo-normal aligned multi-view diffusion model (AN-MVM) for consistent multi-view image generation, which is trained on a set of paired albedo and normal data rendered from 3D object datasets. Then, following the simplified Diesney BRDF model [2], we define a 3D representation that includes albedo, normal, roughness, and metallic properties, which are optimized based on the trained diffusion model and Stable Diffusion [37] model. Compared with previous text-to-3D methods [34, 18, 45, 40, 55], our approach is able to disentangle the illumination and PBR material, achieving high-quality relightable objects under different ambient lighting conditions, as shown in last three columns of Fig.1.
@@ 2.1 Text-to-3D Generation
In recent years, the field of text-to-3D content generation has seen significant advancements, largely inspired by the advancements in text-to-image generation. These advances have particularly been driven by methods employing CLIP-based guidance or score distillation. Methods like  [14, 51, 26] utilize the text-image relationship inherent in vision-language pre-trained models [3, 44], such as CLIP [36] to facilitate general text-to-3D content creation. Another innovative approach, pioneered by DreamFusion [34], employs score distillation to enhance the robustness of text-to-3D generation, which also has been further developed and expanded in subsequent works [18, 45, 55, 42]. The central to both of these paradigms is the use of pre-trained text-to-image diffusion models as a foundation, enabling the creation of diverse and imaginative 3D content. Furthermore, recent advancements have been developed by works [58, 40, 21, 19], which innovatively employ models derived from the Objaverse [7] 3D dataset. These models are used to render 2D RGB images, aiming to train text-to-multi-view image diffusion models. This approach represents a significant stride in optimization-based 3D model generation, as it enables the simultaneous generation of multiple coherent images. Such a technique effectively addresses the challenge of inconsistent directions in geometric optimization, thereby enhancing the consistency and quality of 3D model generation.
In a different paradigm, some methods have shifted towards training diffusion or reconstruction models directly on paired text-3D data. This strategy enables the creation of 3D models that inherently possess text-like semantics. A variety of techniques [6, 59, 13, 30, 15, 9, 10] fall under this category of 3D diffusion generation. In these models, textual information serves as a conditional input, guiding the generation process. This approach emphasizes the manipulation of the underlying 3D data representation, ensuring that the resulting models are both semantically rich and accurate representations of the text descriptions. Moreover, innovative strides have been taken in the realm of 3D reconstruction methods, particularly those grounded in transformer models, exemplified by [12, 17, 50, 46, 53, 47, 56, 43, 61]. These methods introduce a novel perspective, enabling the generation of high-quality 3D models from text or images within seconds, courtesy of their efficient reconstruction networks. The adoption of 3D diffusion and reconstruction methodologies has gained prominence due to their impressive speed in generating 3D objects.
The 2D multi-view diffusion approach and the 3D reconstruction technique utilizing 3D data have provided substantial inspiration. UniDream deeply rethinks the essential principles of these methods and constructed with integrating the strengths of them.
@@ 2.2 Materials Generation
Drawing inspiration from recent material estimation techniques, newer research has focused on generating surface materials for 3D objects. For instance, Fantasia3D [5] combines a physical differential rendering pipeline with SDS to produce detailed 3D objects with realistic surface materials. However, this approach sometimes mixes albedo with reflected light, resulting in blurred material property distinctions. To address this, MATLABER [52] employs a latent BRDF auto-encoder, trained on an extensive dataset, to more effectively differentiate these properties. Building upon this, our method initiates with fixed albedo and normal generation and progressively incorporates other BRDF parameters, achieving a more natural and effective decomposition of surface material of 3D objects.
@@ 3.1 Albedo-Normal Aligned Multi-view Consistent Diffusion Model
Multi-Domain Self-Attention. Based on multi-view consistency and sharing a similar perspective to recent work [22], we further introduce multi-domain consistency. Specifically, we introduce a distinct class label $L$ for the normal domain, and use a two-layer multi-layer perceptron (MLP) to encode this class label $L$ to obtain feature $F_{l}$ with the same dimension with the time-step features $F_{t}$, and $F_{l}$ is added to $F_{t}$ to control the generation process within the normal domain. Subsequently, we apply the self-attention mechanism to the corresponding views between the albedo and normal domains to ensure domain consistency.
@@
Text and Image Semantic Alignment. In order to solve the problem of potential semantic generalization loss caused by only using less 3D synthetic data during AN-MVM training, we use joint training to combine the 2D LAION-Aesthetics data with the published 3D data. Specifically, in line with the approach used by MVDream [40] during our training process, each batch is randomly composed by 3D data or 2D data based on a probability distribution: there’s a 70% chance of using 3D data and a 30% chance of incorporating 2D LAION-Aesthetics data. In this setting, while the normal domain is distinguished by class label, the albedo and 2D LAION-Aesthetics data are not. Moreover, to further differentiate between 3D and 2D data, we add the phrase ", 3D asset" to the captions of 3D data. This strategic inclusion of a significant proportion of 2D data plays a crucial role in enhancing the semantic alignment between the generated image content and the corresponding input text.
@@ 3.2 Transformer-Based Reconstruction Model
Inspired by LRM [12] and Instant3D [17], we have integrated reconstruction models into our text-to-3D generation pipeline, aiming to provide an initial 3D prior for enhancing text-to-3D generation performance. As illustrated in the TRM module of Fig.3, for each object in the AN-MVM training dataset, we randomly select four views $I_{i}$, identical elevation but orthogonal views. Along with these views, the corresponding camera parameters, $C_{i}$, are used as inputs to the model. In this framework, the pre-trained ViT-based DINO-v2 [32] model, denoted as $F$ is employed to extract image features, $F_{i}$, from the albedo images of the four selected views. Concurrently, a learnable camera modulation module processes the camera parameters for each view using $\text{MLP}^{\text{mod}}(C_{i})$ and seamlessly integrates these encoded parameters into the image features. Then we employ learnable tokens, denoted as $T_{h}$, as input to the transformer model. These tokens are designed to undergo cross-attention with the image features, allowing for an effective fusion of the input image information. Subsequently, this is followed by the integration of cross-attention and self-attention modules and a multi-layer perceptron (MLP) to form a transformer block. Multiple such transformer blocks work in sequence to decode the input tokens $T_{h}$ into refined triplane representations $T_{h}^{\prime}$. The representations correspond to the semantic information of the input image. Finally, we decode these triplane representations using a MLP decoder of NeRF to reconstruct the final 3D model, denoted as $\text{Gen}_{\text{3D}}$. This entire process is detailed in Eq.1. And $M$ is the number of transformer layers.
@@
Our approach differs from the configurations used in Instant3D in several key aspects to better adapt our model’s requirements. Firstly, we remove the intrinsic camera parameters of the input multi-view images and only normalize and encode the extrinsic parameters to adapt the multi-view images output by AN-MVM. Furthermore, we use albedo instead of RGB for training to prevent the impact of lighting and shadows in RGB images on the triplane-NeRF reconstruction results. Additionally, in order to increase the resolution of reconstructed model and save training cost, we resize the reference views between $128\times 128$ and $256\times 256$ resolution, and randomly crop $128\times 128$ images from the resized views to supervise the aligned region of rendered images. Finally, in the supervision, we not only render albedo for supervision but also incorporate normal supervision to significantly accelerate the model’s convergence speed and promote the geometry details. These strategic enhancements enable our TRM to deliver superior reconstruction results while reducing the training cost, which will effectively increase the effectiveness of subsequent modules.
@@ 3.4 Physically-Based Rendering (PBR) Material Generation
Different from RichDreamer [35] generating route of PBR material, based on the geometry and albedo from the DMTet refinement, we employ the Stable Diffusion [37] model to generate the PBR material. We adopt the PBR material and illumination representations in Nvdiffrec [28], which is recognized for its speed and efficiency.
@@ 4.1 Implementation Details
Score Distillation Sample(SDS) Refining Details. We implement the refinement stages of NeRF and DMTet based on the Threestudio [^3] https://github.com/threestudio-project/threestudio. Specifically, we render four-view albedo and normal maps at the same time for SDS training. In the NeRF and DMTet refinement stages, we train 5,000 and 2,000 iterations respectively. During training, the loss weights of the albedo and normal domains are 0.8 and 0.2, and an ‘unsharp’ operation [48] is used in the last 500 iterations of each stage.
4.2 Qualitative Comparisons
4.3 Quantitative Evaluations
@@
Figure 5: Comparison of multi-view results generated by MVDream (Row-1) and UniDream (Row-2-3).
Figure 6: Intermediate results of the UniDream generation process. From left to right: the results of TRM reconstruction, the albedo results and the normal results via SDS refinement based on the TRM result using the AN-MVM model, and the PBR result vis SDS generation based on the albedo and normal results using Stable Diffusion model.
@@ 5 Ablation and Analysis
Step-by-Step Comparison of Visualization During 3D Generation. By employing multiple views (as depicted in Fig.5) as input, UniDream’s step-by-step 3D generation results are shown in Fig.6. The left group shows the reconstruction results of TRM. Although the results are relatively rough, they still maintain fairly clear texture and geometric boundaries. The middle two groups present the generated albedo and normal results through SDS refinement using the AN-MVM model, based on the TRM results. After refinement, they exhibit a high quality of 3D models. The right group displays the generated PBR results by SDS optimization using the Stable Diffusion model with fixed albedo and normals, showcasing realistic 3D textures. The quality improves progressively from left to right, reflecting the excellent 3D generation capabilities of UniDream.
@@
Relighting Comparison Driven by Environment Maps. Fig.8 shows the different environment maps used by UniDream. Fig.8(a) is the environment map used by Fig.7 and the visualize of UniDream’s overall capabilities as shown in Fig.4. Fig.8(b) and Fig.8(c) are the environment maps used by ’Relighting-I’ and ’Relighting-II’ in the teaser of UniDream respectively. Changing different environment maps will produce different rendering results, which reflects UniDream’s excellent relightability.